Company A has overstated the goodwill of its acquisition of Company B in 2018. The price of acquisition was 257. Company A estimated fair value of net assets to be 103. In 2019, you as an analyst want to make an accounting adjustment to the recorded goodwill by impairing 50% of it based on newly acquired information. With what amount will the net profit/loss (after tax) item on the income statement 2019 decrease based on this adjustment? Consider that the tax rate is 20%. Give your answer in absolute terms (not with minus sign).
Answer #1

Goodwill = Price of acquisition - Fair Value of Net Assets

Goodwill originally recognized by company A on acquisition = 257-103 = 154

Impairment in Goodwill = 154 * 50% = 77

Tax benefit due to impairment = 77 * 20% = 15.40

Thus, Net decrease in the Net Profit (after tax) = (77 -15.40) = 61.60